Completely factorize the following polynomial: x4−3x3+4x2−8


1
Expert's answer
2021-11-29T19:35:50-0500
"x^4-3x^3+4x^2-8"

"=(x^4-2x^3)-(x^3-2x^2)+(2x^2-8)"

"=x^3(x-2)-x^2(x-2)+2(x-2)(x+2)"

"=(x-2)(x^3-x^2+2x+4)"

"=(x-2)((x^3+x^2)-(2x^2+2x)+(4x+4))"

"=(x-2)(x+1)(x^2-2x+4)"

"=(x-2)(x+1)(x-2)^2"

"=(x-2)^3(x+1)"

"x^4-3x^3+4x^2-8=(x+1)(x-2)^3"